Raymond’s Furniture Outlet budgets $30,000 for advertisement costs promoting their furniture sales. Television ads cost $500 each and radio ads cost $100 each. The company wants to buy a total of 220 ads. a) Give the system of equations for solving this situation. b) Find the number of television ads and radio ads that the company should buy.


Please help me answer this question!!

Answer by Annabelle1(69)

let x= the number of tv advertisements and y = the number of radio advertisements.
x+y=220 .........................1
sincew we know that the tv ads cost $500 and the radio ads $100 the total costs will be given by
500x+100y=30000..................2
we multiply the first equation by 100 and we have
100x+100y=22000..................3
equation 2 - 3
(500x-100x)+(100y-100y)=30000-22000
400x=8000
x=20
x+y=220
20+y=220
y=200
20 tv ads and 200 radio